OpenStudy (anonymous):

If the bases are 1/2 and 1/4, and the height is 1, what is the area? 1/8 sq. units 3/8 sq. units 3/4 sq. units

OpenStudy (anonymous):

a=l*w

OpenStudy (anonymous):

A=LxWxH A=1/2x1/4x1 A=1/8 sq units

OpenStudy (anonymous):

no need for the height your now computing for the volume

OpenStudy (anonymous):

so 1/2x1/4=1/8

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wierd question came with height

OpenStudy (anonymous):

we can use the height if the question is finding the volume or surface area...

OpenStudy (anonymous):

true, but question posted didn't say if area of what also whether area or surface area